A common arrangement in real estate lending might call for a 5 year loan with, say, a 15 years amortization. What this means is that the borrower makes a payment every month of a fixed amount based on a 15 year amortization. However, after 60 months, the borrower makes a single, much larger payment called a “balloon’ or “bullet” to pay off the loan. Because the monthly payments don’t fully pay off the loan, the loan is said to be partially amortized.

Suppose we have a $100,000 commercial mortgage with a 12 percent APR and a 20 year (240 month) amortization. Further suppose the mortgage has a five year balloon.

1)What will the monthly payment be?

2)How big will the balloon payment be?
Answer #1

1: Using financial calculator
Input: PV = -100000, N=240, I/Y=12/12=1

Solve for PMT as 1101.09

2: Balloon payment =

91744.33

Workings

1 Monthly payment $1,101.09
2 Amount repaid in 5 years -8255.67
Balance loan 91744.33

Book1- Excel AutoSave Oi Sign in Comments Share O Tell me what you want to do File Home Formulas Review View Help Insert Draw